OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (1, –7); Slope: -2/3 (negative 2 over three)

OpenStudy (anonymous):

A. y-7=2/3(x-1) B.y+7=2/3(x+1) C.y-7=-2/3(x+1) D.y+7=-2/3(x-1)